Title: Euler–Poisson systems as energy minimizing paths in the Wasserstein space P_{2}(ℝ) Abstract: This paper uses a variational approach to establish existence and uniqueness

of solutions (σ_{t}, v_{t}) of the 1–d Euler-Poisson system, mimimizing an action. Here, σ_{0}, σ_{T} are prescribed and the time interval [0, T] satisfies T < π.This extends the concept of the Euler-Poisson system to general measures. These solutions conserve the Hamiltonian, Legendre transform of the Lagrangian appearing in the action. They yield a path t → σ_{t} on P the set of probability measures on the real line and turn out to be characteristic of an infinite dimensional Hamilton-Jacobi equation on P. The associated Hamiltonian system is nothing but the 1–d Vlasov-Poisson system. When σ_{t} = δ_{y(t)} is a dirac mass, the Euler-Poisson system reduces to ÿ + y = 0, the Hamilton-Jacobi equation is merely finite dimensional and is given by ∂_{t}u + 1/2(∂_{y}u)^{2} + 1/2y^{2} + 1/24 = 0. (This is a joint work with T. Nguyen and A. Tudorascu).